OpenStudy (anonymous):

A cart is moving horizontally along a straight line with constant speed 30m/s. A projectile is to be fired from the moving cart in such a way that it will return to the cart after the cart has moved 80m . At what speed (relative to cartt) must the projectile launched? g=10

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

u were offline at that time!

OpenStudy (anonymous):

i posted this once but i didnt understand

OpenStudy (anonymous):

The time taken by the cart to travel 80m = 8/3 sec. This is equal to the time of flight of the projectile = 2u sinx/g = 8/3, where u is the initial velocity and x is the angle at which the projectile is projected. u*sin x= 40/3 . sin x = 40/3u , cos x = sqrt( 1-sin^2 x ) = (sqrt( 9u^2-1600))/3u . The horizontal range of the projectile is given by u^2*sin 2x/g=80m . u^2*sin 2x = 800 , sin 2x = 2*sin x*cos x = 40(sqrt( 9u^2 - 1600 ))/9u^2 Plugging this in the range equation , we get 40(sqrt( 9u^2 - 1600 ))/9=800, sqrt(9u^2-1600) = 180 , squaring on both sides and simplifying , 9u^2=30800 , u= 58.5 m/sec.

OpenStudy (anonymous):

My options are 10 10sqrt8 40/3 none of these

OpenStudy (anonymous):

I think it is none of these.Not sure.Rechecking the values now.Please wait.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think u r correct

OpenStudy (anonymous):

Oops I made a mistake, here is the new solution: The time taken by the cart to travel 80m = 8/3 sec. This is equal to the time of flight of the projectile = 2u sinx/g = 8/3, where u is the initial velocity and x is the angle at which the projectile is projected. u*sin x= 40/3 . sin x = 40/3u , cos x = sqrt( 1-sin^2 x ) = (sqrt( 9u^2-1600))/3u . The horizontal range of the projectile is given by u^2*sin 2x/g=80m . u^2*sin 2x = 800 , sin 2x = 2*sin x*cos x = 40(sqrt( 9u^2 - 1600 ))/9u^2 Plugging this in the range equation , we get 80(sqrt( 9u^2 - 1600 ))/9=800, sqrt(9u^2-1600) = 90 , squaring on both sides and simplifying , 9u^2= , u= 32,829 m/sec. Still getting none of the above option.I've checked my calculation.

OpenStudy (anonymous):

wat .......32,829 m/sec.

OpenStudy (anonymous):

man..........its huge

OpenStudy (anonymous):

i think we have to consider relative motion

OpenStudy (anonymous):

as it is thrown from the moving cart

OpenStudy (anonymous):

Sorry it is supposed to be a point.It is 32.829 m/sec

OpenStudy (anonymous):

And yes it is relative to the velocity of the cart.

OpenStudy (anonymous):

ok............thxxx